# Interpretable LLM-based Table Question Answering

### v2

@@ D.3. An example of data-attribution tracking for Table QA
Appendix E Hallucinations in Sub-table Selection
@@
As illustrated in Figure 12, although Chain-of-Table correctly answers the question Q: True or False? In four different baseball games, the final score was 9-2, it irrationally selects unrelated information (game 3) from the input table. Similarly, DATER, shown in Figure 13, selects rows 2, 3, 4, 5, and 9 to answer the same question. However, the inclusion of row 3 is illogical and does not contribute to a valid answer.
Figure 12. Although CoTable correctly answers the question Q: True or False? In four different baseball games, the final score was 9-2, it irrationally selects unrelated information (game 3) from the input table.
Figure 13. DATER selects rows 2, 3, 4, 5, and 9 to answer the question. However, the inclusion of row 3 is illogical and does not contribute to a valid answer.
Appendix F Details for LLM-as-a-Judge Experiments
F.1. Prompt to LLM Judges in Forward Simulation
@@
Based on this explanation, guess what the model will predict on the Statement based on the provided explanation. Answer with only ‘TRUE’ or ‘FALSE’: """
F.2. Prompt to LLM Judges in Preference Ranking
@@
"Text2SQL": """Text2SQL is a method that translates the natural language query into a single SQL query. The SQL query itself serves as the explanation for how the system arrives at its answer. The explanation contains: The generated SQL command that will be directly applied onto the table to generate the final answer.""",
@@
Appendix G Prompt Engineering
G.1. Prompt for Atomic Planning
G.1.1. Decomposition of Query $Q$
The decomposition process breaks down the complex query $Q$ into a sequence of atomic steps. This is achieved through a carefully crafted prompt provided to the LLM. The prompt includes:
@@ G.1.1. Decomposition of Query $Q$
  • In-context Examples: We provide example inputs $(T,Q)$ along with their corresponding plans to serve as in-context examples for planning (see Appendix H).
G.1.2. Sequencing of Steps
@@
G.1.3. The Importance of the Step Order
@@ G.2. Prompt for Step-to-SQL
Appendix H In-context Examples
H.1. In-context Examples for Atomic Planning
@@
H.2. In-context Examples for Step-to-SQL
@@
Appendix I Error Analysis of POS and Improved Planning Algorithm
We notice that many errors in POS are due to the planning stage rather than Step-to-SQL process. The Planner often misses condition checks (see Figure 14Figure 16Figure 15Figure 17Figure 18) in atomic steps. An interesting (and inherently unavoidable) error due to the exact-matching nature of SQL can also be found in Figure 19.
Based on this observation, we implement an improved planning algorithm in which only one step is generated at a time, rather than generating all steps upfront, as shown in Figure 1. This approach encourages the LLM to think one step at a time and reduces the complexity of the planning task. The input of NL Planner is the previous steps and the current intermediate table.
As shown in Table 11, planning one step at a time leads to substantial improvements in accuracy for both TabFact and WikiTQ with GPT-4o-mini. Please note that this one-step planning should not change the interpretability of POS. For the lack of proper baselines for Table QA using GPT-4o-mini, we do not put this result into the main text, but the future works should consider using one-step planning over one-time planning.
Table 11. Table QA accuracy (%) in TabFact and WikiTQ using planning one step at a time (one-step planning) with GPT-4o-mini.
Method TabFact WikiTQ
@@ Appendix I Error Analysis of POS and Improved Planning Algorithm
Figure 14. POS (ours) predicts TRUE but the groundtruth is FALSE (False Positive). In planning, POS misses checking the player name.
@@
Figure 15. POS (ours) predicts TRUE but the groundtruth is FALSE (False Positive). In planning, POS misses checking the player name.
@@
Figure 16. POS (ours) predicts TRUE but the groundtruth is FALSE (False Positive). In planning, POS misses checking the score.
@@
Figure 17. POS (ours) predicts FALSE but the groundtruth is TRUE (False Negative). In planning, POS misses checking the year.
@@
Figure 18. POS (ours) predicts FALSE but the groundtruth is TRUE (False Negative). In planning, POS should select two rows at the same step.
Figure 19. POS (ours) predicts FALSE but the groundtruth is TRUE (False Negative). The exact-matching nature of SQL makes POS cannot retrieve the relevant information.
Appendix J Human Study Interface
To evaluate the effectiveness of different explanation methods in Table QA on human users, we develop a web-based interface using HuggingFace Gradio and Flask. The interface is designed for Forward Simulation to guide participants through the study seamlessly, ensuring they understand the tasks and provide valuable feedback.
Overview of the Forward Simulation Interface Flow:
  1. (1)
    Informed Consent $\Rightarrow$
  2. (2)
    Introduction to Table QA and Forward Simulation $\Rightarrow$
  3. (3)
    Introduction to Explanations in Table QA $\Rightarrow$
  4. (4)
    Welcome page where users are asked to choose one of 4 XAI methods $\Rightarrow$
  5. (5)
    Specific explanation page for the chosen method $\Rightarrow$
  6. (6)
    Experiment pages for 10 samples $\Rightarrow$
  7. (7)
    Completion page!
J.1. Informed Consent
Informed Consent for Table QA Study Study Information: You are invited to participate in a research study on Table QA systems. This study aims to improve how AI systems explain their reasoning when answering questions based on tabular data. Study Duration and Requirements: 1. The entire study will take approximately 20 minutes to complete. 2. Please perform this study on a computer (not a phone). 3. Do not seek help from the Internet or other people during the study. Study Structure: 1. Introduction to Table QA and task explanation 2. Main study: 10 questions about Table QA explanations Benefits: Your participation will contribute to the development of AI systems that can better explain their reasoning to humans, particularly in the domain of question answering from tabular data. There are no known risks associated with this study. Data Usage and Confidentiality: All data collected will be anonymized and used solely for research purposes. Your personal information will be kept confidential. Voluntary Participation: Your participation in this study is entirely voluntary. You may choose to withdraw at any time without any consequences. Contact Information: If you have any questions or concerns about this study, please contact [anonymized]. Agreement: By clicking “I Agree” below, you confirm that you have read and understood this informed consent, and you agree to participate in this Table QA study under the terms described above. I Agree
J.2. Introduction to Table QA and Forward Simulation
Introduction to Table QA In this experiment, you will interact with Table QA models. Table QA involves answering questions based on data provided in tables. Verify if the following Statement is TRUE or FALSE Statement: The Wildcats kept the opposing team scoreless in four games. Input Table Caption: 1947 Kentucky Wildcats Football Team Game Date Opponent Result Wildcats Points Opponents Record 1 $9$-09-20 Ole Miss Loss 7 14 0 - 1 2 $9$-09-27 Cincinnati Win 20 0 1 - 1 3 $9$-10-04 Xavier Win 20 7 2 - 1 4 $9$-10-11 9 Georgia Win 26 0 3 - 1, 20 5 $9$-10-18 10 Vanderbilt Win 14 0 4 - 1, 14 6 $9$-10-25 Michigan State Win 7 6 5 - 1, 13 7 $9$-11-01 18 Alabama Loss 0 13 5 - 2 8 $9$-11-08 West Virginia Win 15 6 6 - 2 9 $9$-11-15 Evansville Win 36 0 7 - 2 10 $9$-11-22 Tennessee Loss 6 13 7 - 3 Model thinks this Statement is: TRUE Model Simulation Task Given an input statement, an Artificial Intelligence (AI) model will output either TRUE or FALSE. Your job in this Simulation task is to use the AI’s explanation to guess the machine response. Specifically, please choose which response (Statement is TRUE/Statement is FALSE) the model would output regardless of whether you think that response is correct or not. Next
J.3. Introduction to Explanations in Table QA
Understanding Attribution Explanations Attribution explanations highlight specific parts of a table—such as rows, columns, or cells—that are most relevant to the answer predicted by a Table QA model. These explanations help you understand which information of the input the system considered important when predicting the answer. Table caption: 1947 Kentucky Wildcats Football Team Statement to verify: “The Wildcats kept the opposing team scoreless in 4 games.” Game Date Opponent Result Wildcats Points Opponents Record 1 $9$-09-20 Ole Miss Loss 7 14 0 - 1 2 $9$-09-27 Cincinnati Win 20 0 1 - 1 4 $9$-10-11 9 Georgia Win 26 0 3 - 1 , 20 5 $9$-10-18 10 Vanderbilt Win 14 0 4 - 1 , 14 9 $9$-11-15 Evansville Win 36 0 7 - 2 In this example, the Table QA model has highlighted specific rows and cells to explain its prediction: 1. The entire rows for games 2, 4, 5, and 9 are highlighted in yellow. 2. Within these rows, the Opponents column cells containing “0” or “scoreless” are highlighted in green. These highlights indicate that the system identified four games where the opposing team did not score, verifying the statement as TRUE. The yellow highlighting shows the relevant rows, while the green highlighting represents the cells containing fine-grained information needed to verify the statement. By using different colors for highlighting, the system provides a more nuanced explanation: 1. Yellow highlights (rows): Show the overall context of the relevant games. 2. Green highlights (cells): Pinpoint the exact information (opposing team’s score of 0) that directly answer the question. During the experiment, you will use explanations to choose which response (Statement is TRUE/ Statement is FALSE) the model would output, regardless of whether you think that response is correct or not. Proceed to Experiment
J.4. Welcome Page
[…]
J.5. Experiment Page
Sample: 1 / 10 Please note that in select row function, starting index is 0 for Chain-of-Table and 1 for DATER and Index * represents the selection for all rows. Based on the explanation below, please guess what the AI model will predict on the input Statement below. [Input Statement] [Explanation content] Guess what the model will predict on the Statement based on the provided explanation? Model will predict: Statement is TRUE Model will predict: Statement is FALSE
J.6. Completion Page
Thank you! You’ve successfully completed the experiment. Your predictions have been recorded. Your Labeling Accuracy user_accuracy % You Predicted TRUE true_percentage % You Predicted FALSE false_percentage % Back to Start Page

### v3

@@ E.3 An example of data-attribution tracking for Table QA
Appendix F Hallucinations in Sub-table Selection
@@
As illustrated in Fig. 12, although Chain-of-Table correctly answers the question Q: True or False? In four different baseball games, the final score was 9-2, it irrationally selects unrelated information (game 3) from the input table. Similarly, DATER, shown in Fig. 13, selects rows 2, 3, 4, 5, and 9 to answer the same question. However, the inclusion of row 3 is illogical and does not contribute to a valid answer.
Figure 12: Although CoTable correctly answers the question Q: True or False? In four different baseball games, the final score was 9-2, it irrationally selects unrelated information (game 3) from the input table.
Figure 13: DATER selects rows 2, 3, 4, 5, and 9 to answer the question. However, the inclusion of row 3 is illogical and does not contribute to a valid answer.
Appendix G Details for LLM-as-a-Judge Experiments
In this section, we describe how we prompt LLMs to work as judges for evaluating explanation methods.
G.1 Prompt to LLM Judges in Prediction Verification
@@
The Table Fact Verification (TabFact) model is working on verifying if a given Statement is TRUE or FALSE based on a given input Table.
You are given an HTML file containing a Statement, Input Table, Prediction, and an Explanation clarifying the Prediction.
Your task is to carefully analyze the Explanation and determine whether the Prediction is correct or not.
Explanation Method: [method]
[method_specific_info]
HTML content: [html_content]
Answer with ’option1’ or ’option2’ only.
You MUST ignore the order of the options and answer based on the correctness of the Prediction!
G.2 Prompt to LLM Judges in Forward Simulation
{mdframed}
[…]
Given an input statement, an Artificial Intelligence (AI) model will output either TRUE or FALSE. Your job in this Simulation task is to use the AI’s explanation to guess the machine response. Specifically, please choose which response (TRUE/FALSE) model would output regardless of whether you think that response is correct or not.
@@ G.2 Prompt to LLM Judges in Forward Simulation
Based on this explanation, guess what the model will predict on the Statement based on the provided explanation. Answer with only ‘TRUE’ or ‘FALSE’:
G.3 Prompt to LLM Judges in Preference Ranking
@@
"Text2SQL": """Text2SQL is a method that translates the natural language question into a single SQL query. The SQL query itself serves as the explanation for how the system arrives at its answer. The explanation contains: The generated SQL command that will be directly applied onto the table to generate the final answer.""",
@@
Appendix H Prompt Engineering
H.1 Prompt for Atomic Planning
H.1.1 Decomposition of Question $Q$
The decomposition process breaks down the complex question $Q$ into a sequence of atomic steps. This is achieved through a carefully crafted prompt provided to the LLM. The prompt includes:
@@ H.1.1 Decomposition of Question $Q$
  • In-context Examples: We provide example inputs $(T,Q)$ along with their corresponding plans to serve as in-context examples for planning (see Appendix I).
H.1.2 Sequencing of Steps
@@
H.1.3 The Importance of the Step Order
@@ H.2 Prompt for Step-to-SQL
Appendix I In-context Examples
I.1 In-context Examples for Atomic Planning
@@
I.2 In-context Examples for Step-to-SQL
@@
Appendix J Error Analysis of POS and Improved Planning Algorithm
J.1 Error Analysis of POS
We notice that many errors in POS are due to the planning stage rather than the Step-to-SQL process. In particular, the Planner misses condition checks (see Fig. 14Fig. 16Fig. 15Fig. 17Fig. 18) in atomic steps. Another interesting (and inherently unavoidable) error is presented in Fig. 19. Arguably, this error can be attributed to the planner, as it has access to both the query and the input table and should therefore generate the correct step. In this case, the Step-to-SQL component is functioning as intended. We also offer an alternative perspective on this interesting failure in Sec. J.2.
J.2 Quantifying contributions of Planning and Step-to-SQL to POS errors
In the example shown in Fig. 19, the failure to retrieve the correct rows is not due to any flaw in our Step-to-SQL converter. Instead, it stems from a mismatch between the query token and the table’s content—and, ultimately, from an inconsistent ground-truth label. Specifically, the NL Planner faithfully mirrors the input string “bjørn” when it generates its first step (“Select rows where player = ‘thomas bjørn’”), but the actual table only contains “thomas bjarn,” so no rows are returned. Under standard Table-QA semantics, “bjørn” and “bjarn” denote two distinct entities, so the correct label for this query should be False, not True. Consequently, neither the Step-to-SQL module nor the NL Planner has malfunctioned; the discrepancy arises from the dataset’s labeling and the lack of accent normalization in the query definition.
Putting this example in context, we observe that across all of our error analyses there are no instances where the Step-to-SQL module itself directly produces an incorrect SQL query. Instead, errors occur at two other points in the pipeline:
  • NL Planner errors: the natural-language plan deviates from the correct reasoning (even though every resulting SQL statement remains executable), or
  • Fallback errors: at least one SQL in the plan is invalid, causing the system to defer to the end-to-end QA model, which then also fails to produce the right answer.
To quantify how much each of these failure modes contributes to our overall error rate on TabFact and WikiTQ, we labeled every incorrect sample in our evaluation set as either a Planning error or a Fallback error. The breakdown of their relative frequencies is reported in Tab. 12.
Table 12: Contribution to POS errors of Planning and Fallback.
Dataset Model Planning (%) Fallback (%)
TabFact GPT-4o-mini 95.4 4.6
TabFact GPT-3.5 89.7 10.3
WikiTQ GPT-4o-mini 75.5 24.5
WikiTQ GPT-3.5 70.7 29.3
We observe that planning mistakes constitute the vast majority of errors: over 95% for GPT-4o-mini on TabFact (and nearly 90% for GPT-3.5), and around 70–75% on the more complex WikiTQ benchmark. This confirms that planning is the primary bottleneck in POS.
J.3 Planning one step at at time improves Table QA accuracy
Based on this observation, we implement an improved planning algorithm in which only one step is generated at a time, rather than generating all steps upfront, as shown in Fig. 1. This approach encourages the LLM to think one step at a time and reduces the complexity of the planning task. The input of NL Planner is the previous steps and the current intermediate table.
As shown in Tab. 13, planning one step at a time leads to substantial improvements in accuracy for both TabFact and WikiTQ with GPT-4o-mini. There are leading hypotheses for these improvements:
  • Grounding in real data. Before generating each new step, one-step planning provides the actual intermediate table (the result of the previous SQL) into the prompt. This lets the LLM see exactly which rows and values it’s operating on (condition on the actual intermediate table state), mitigating hallucinations and mis-alignments. By contrast, generating an entire multi-step plan up front requires the model to imagine how each intermediate table will look—and if those imagined tables don’t match reality, later steps can go wrong.
  • Focused context window. By feeding the LLM only the rows and columns it needs for each step—rather than the entire table—we keep the input prompt concise and free of irrelevant data. Since LLMs can struggle when their context grows too large Li et al. (2024), this focused context helps them generate the next step more accurately.
Table 13: Planning one step at a time (one-step planning) with GPT-4o-mini.
Method TabFact (%) WikiTQ (%)
@@
Figure 14: POS predicts TRUE but the groundtruth is FALSE (False Positive). In planning, POS misses checking the player name.
@@
Figure 15: POS predicts TRUE but the groundtruth is FALSE (False Positive). In planning, POS misses checking the player name.
@@
Figure 16: POS predicts TRUE but the groundtruth is FALSE (False Positive). In planning, POS misses checking the score.
@@
Figure 17: POS predicts FALSE but the groundtruth is TRUE (False Negative). In planning, POS misses checking the year.
@@
Figure 18: POS predicts FALSE but the groundtruth is TRUE (False Negative). In planning, POS should select two rows at the same step.
Figure 19: POS predicts FALSE but the groundtruth is TRUE (False Negative). Arguably, this error can be attributed to the planner, as it sees both the query and the input table and should therefore generate the correct step (i.e., Select rows where ‘player’ is ‘thomas bjarn’). We observe that the Step-to-SQL component always functions correctly. We also offer an alternative perspective on this interesting failure in Sec. J.2.